### v1

# Review of Degenerate Higher Order Scalar Tensor Theories in Cosmology

Abstract
Degenerate Higher Order Scalar Tensor (DHOST) theories are the most general scalar-tensor theories whose Lagrangian depends on the metric tensor and a single scalar field and which propagate only one scalar degree of freedom, without being plagued by Ostrogradsky instabilities. This is achieved through certain degeneracies of the functions forming their Lagrangian. They generalise the Horndeski and beyond-Horndeski theories. Originally build to describe the late-time acceleration of the expansion of the universe, generalising the cosmological constant, they can be used to build models for early universe, to describe inflation or alternatives to standard inflation. In the late universe, they modify the standard Vainstein screening mechanism from Horndeski theories which can have observable consequences and are suited to build black hole models, featuring non-stealth Kerr black hole solutions. In this work we review their effects on cosmology, looking at their basic properties, their parameterisations and classifications, and then focusing on solutions in the early and the late universe.
1 Introduction
The General Theory of Relativity, proposed more than a century ago, supplemented by a cosmological constant, has been successfully used to describe the universe. Its predictions are in agreement with the observations, including the accelerated expansion of the universe and the presence of gravitational waves.
The nature of the cosmological constant is, however, unexplained and this has led to a large number of theories being developed, trying to relate this to the unknown nature of the dark energy, but also to find a way to parameterise it and to constrain it with various cosmological probes. Moreover, extensions of the theory (modified gravity) are required to account for the presence of singularities such as black holes.
In the very early universe, just after in the initial Big bang, there was a brief period of exponential expansion, the inflationary epoch, developed from the late 1970s by several theoretical physicists including Alexei Starobinsky, Alan Guth, Stephen Hawking and Andrei Linde. This theory provides a satisfactory explanation for the presence of the structures that we observe in the universe today through the rapid growth of primordial quantum fluctuations [1, 2, 3, 4].
As these two early and late universe epochs are qualitatively similar, theories that are developed for one era can be adapted to the other. A classical method to parameterise such theories is to introduce scalar and tensor fields to model the interactions. In this work we concentrate on theories involving one single scalar field. It was long thought that the most general such theories that do not yield additional degrees of freedom and hence avoid the generation of Ostrogradsky ghost instabilities [5] need to have second-order Euler-Lagrange equations. These are the Horndeski theories [6]. It turns out that this last requirement is not necessary, and they have been generalised to beyond Horndeski [7, 8, 9] and finally to the Degenerate Higher Order Scalar Tensor (DHOST) theories [10]. The latter include all the scalar-tensor theories having these properties, including Horndeski and beyond Horndeski ones.
The early developments of DHOST theories are reviewed in [11], where the consequences of these theories in the late universe and on dark energy are discussed. The paper presents the emergence of DHOST theories through the generalisation of Horndeski and beyond Horndeski ones and discusses their cosmological consequences as well as the gravitational properties of astrophysical bodies in this context. The mechanisms concerning the degeneracies are presented in detail, followed by a classification of the models and then the study of the theories within the formalism of the Effective Field Theory of dark energy and screening mechanisms.
In this review, we discuss the uses of DHOST theories in cosmology in both phases of accelerated expansion of the universe. We introduce the DHOST theories in Section 2, and briefly mention the traditional classification of quadratic DHOST theories. The Effective Field Theory of Dark Energy and its implications on DHOST models is then presented in Section 3. The consequences of DHOST theories on astrophysical bodies (neutron stars and black holes) and the Vainshtein mechanism are the subject of Section 4. The DHOST model for inflation is presented in Section 5, while bouncing cosmology with DHOST models is reviewed in Section 6. We conclude in Section 7.
2 DHOST Theories
DHOST theories involve derivatives up to second order in the scalar field $\phi$ and can be classified in terms of the powers of derivatives present. So far, quadratic [10, 12, 13, 14] and cubic [15] theories have been studied. The most general action containing up to cubic terms is given by
$S[g,\phi]$ $=\int d^{4}x\,\sqrt{-g}\left[F_{(2)}(X,\phi)\,{}^{(4)}R+P(X,\phi)+Q(X,\phi)\square\phi+\sum_{i=1}^{5}A_{i}(X,\phi)L^{(2)}_{i}\right.$
$+\left.F_{(3)}(X,\phi)\,G_{\mu\nu}\phi^{\mu\nu}+\sum_{i=1}^{10}B_{i}(X,\phi)L^{(3)}_{i}\right]\,,$ (2.1)
where
$\phi_{\mu}\equiv\nabla_{\mu}\phi\,,\qquad\phi_{\mu\nu}\equiv\nabla_{\nu}\nabla_{\mu}\phi\,,\qquad X\equiv\phi^{\mu}\phi_{\mu}\,$ (2.2)
and the elementary quadratic and cubic Lagrangians are
$L^{(2)}_{1}=\phi_{\mu\nu}\phi^{\mu\nu}\,,\qquad L^{(2)}_{2}=(\Box\phi)^{2}\,,\qquad L^{(2)}_{3}=(\Box\phi)\phi^{\mu}\phi_{\mu\nu}\phi^{\nu}\,,$ (2.3)
$L^{(2)}_{4}=\phi^{\mu}\phi_{\mu\rho}\phi^{\rho\nu}\phi_{\nu}\,,\qquad L^{(2)}_{5}=(\phi^{\mu}\phi_{\mu\nu}\phi^{\nu})^{2}\,,$ (2.4)
and
$L^{(3)}_{1}=(\Box\phi)^{3}\,,\quad L^{(3)}_{2}=(\Box\phi)\,\phi_{\mu\nu}\phi^{\mu\nu}\,,\quad L^{(3)}_{3}=\phi_{\mu\nu}\phi^{\nu\rho}\phi^{\mu}_{\rho}\,,$ (2.5)
$L^{(3)}_{4}=\left(\Box\phi\right)^{2}\phi_{\mu}\phi^{\mu\nu}\phi_{\nu}\,,\quad L^{(3)}_{5}=\Box\phi\,\phi_{\mu}\phi^{\mu\nu}\phi_{\nu\rho}\phi^{\rho}\,,\quad L^{(3)}_{6}=\phi_{\mu\nu}\phi^{\mu\nu}\phi_{\rho}\phi^{\rho\sigma}\phi_{\sigma}\,,\qquad$ (2.6)
$L^{(3)}_{7}=\phi_{\mu}\phi^{\mu\nu}\phi_{\nu\rho}\phi^{\rho\sigma}\phi_{\sigma}\,,\quad L^{(3)}_{8}=\phi_{\mu}\phi^{\mu\nu}\phi_{\nu\rho}\phi^{\rho}\,\phi_{\sigma}\phi^{\sigma\lambda}\phi_{\lambda}\,,\qquad$ (2.7)
$L^{(3)}_{9}=\Box\phi\left(\phi_{\mu}\phi^{\mu\nu}\phi_{\nu}\right)^{2}\,,\quad L^{(3)}_{10}=\left(\phi_{\mu}\phi^{\mu\nu}\phi_{\nu}\right)^{3}\,.$ (2.8)
These are all the possible contractions of the second-order derivatives $\phi_{\mu\nu}$ with the metric $g_{\mu\nu}$ and with the scalar field gradient $\phi_{\mu}$.
The functions $A_{i}$ and $B_{i}$ must satisfy, however, certain degeneracy conditions such that the theory produces only a single scalar degree of freedom and no ghosts [10, 16, 17, 18].
We note that Horndeski theories are obtained by setting
$F_{(2)}=G_{4}\,,\qquad A_{1}=-A_{2}=2G{4,X}\,,\qquad A_{3}=A_{4}=A_{5}=0\,,$
$F_{(3)}=G_{5}\,,\qquad 3B_{1}=-B_{2}=\frac{3}{2}B_{3}=G_{5,X}\,,\qquad B_{i}=0\ (i=4,\dots,10)\,.$ (2.9)
The degeneracies arising in DHOST theories can be understood as follows. When considering beyond Horndeski and DHOST theories, the corresponding Euler-Lagrange equations can be of higher than second order. They can be equivalently reduced to a system of second-order equations by introducing additional constraint variables. Through a Hamiltonian formulation of this system, one can count the number of degrees of freedom of the system. If the matrix of coefficients is singular, however, then the number of degrees of freedom is reduced. This can be translated into conditions on the functions appearing in the Lagrangian and it can be used to classify the DHOST theories. An intuitive explanation on the degeneracies appearing in DHOST theories is presented in Ref. [10].
The degeneracy conditions place additional constraints on the functions $A_{i}$ and $B_{i}$. Based on whether the theories are quadratic, cubic or mixed, DHOST theories can be classified in several classes: quadratic theories (containing 7 subclasses), cubic theories (with 9 subclasses) and mixed quadratic and cubic theories (25 subclasses). The quadratic theories are presented in detail in [10, 13, 12], and the cubic ones in [15]. In the case of quadratic theories, a geometric interpretation of the degeneracies is discussed in Ref. [19].
As these results are very technical, I shall present in the following only the classification of quadratic theories, leaving the interested reader to discover the classification of the cubic and mixed quadratic/cubic theories in [15].
For quadratic theories, the degeneracy conditions are given by
$D_{0}(X)=0,\qquad D_{1}(X)=0,\qquad D_{2}(X)=0\,,$ (2.10)
where
$D_{0}(X)$ $=-4(A_{1}+A_{2})\left[XF(2A_{2}+XA_{4}+4F_{X})-2F^{2}-8X^{2}F_{X}^{2}\right]\,,$ (2.11)
$D_{1}(X)$ $=4\left[X^{2}A_{2}(3A_{1}+A_{2})-2F^{2}-4XFA_{1}\right]A_{4}+4X^{2}F(A_{1}+A_{2})A_{5}$
$+8XA_{1}^{3}-4(F+4XF_{X}-6XA_{1})A_{2}^{2}-16(F+5XF_{X})A_{1}A_{2}$
$+4X(3F-4XF_{X})A_{2}A_{3}-X^{2}FA_{2}^{2}+32F_{X}(F+2XF_{X})A_{1}$
$-16FF_{X}A_{2}-8F(F-XF_{X})A_{3}+48FF_{X}^{2}\,,$ (2.12)
$D_{2}(X)$ $=4\left[2F^{2}+4XFA_{1}-X^{2}A_{2}(3A_{1}+A-2)\right]A_{5}+4(2A_{1}-XA_{3}-4F_{X})A_{2}^{2}$
$+4A_{2}^{3}+3X^{2}A_{2}A_{3}^{2}-4XFA_{3}^{2}+8(F+XF_{X})A_{2}A_{3}-32F_{X}A_{1}A_{2}$
$+16F_{X}^{2}A_{2}+32F_{X}^{2}A_{1}-16FF_{X}A_{3}\,.$ (2.13)
Here we have dropped the subscript from the function $F_{(2)}$, as we are only dealing with quadratic theories. The classification of the theories is given in terms of how these three equations can be satisfied simultaneously. We express in square brackets the naming used in Ref. [12]
Class I ($A_{1}+A_{2}=0$)
This class is characterised by
$A_{2}=-A_{1}\,.$ (2.14)
This class has two sub-classes,
  • Class Ia [N-I] ($F-XA_{1}\neq 0$)
    There are three free functions, $F$, $A_{1}$ and $A_{3}$. The others are given by:
    $A_{4}$ $=\frac{1}{8(F-A_{1}X)^{2}}\left[4F\left(3(A_{1}-2F_{X})^{2}-2A_{3}F\right)-A_{3}X^{2}(16A_{1}F_{X}+A_{3}F)\right.$
    $\left.\qquad\qquad\qquad+4X\left(3A_{1}A_{3}F+16A_{1}^{2}F_{X}-16A_{1}F_{X}^{2}-4A_{1}^{3}+2A_{3}FF_{X}\right)\right]\,,$ (2.15)
    $A_{5}$ $=\frac{1}{8(F-A_{1}X)^{2}}(2A_{1}-A_{3}X-4F_{X})\left[A_{1}(2A_{1}+3A_{3}X-4F_{X})-4A_{3}F\right]\,.$ (2.16)
  • Class Ib [N-II] ($F-XA_{1}=0$)
    There are three free functions, $F$, $A_{4}$ and $A_{5}$ and
    $A_{3}=\frac{2\left(F-2XF_{X}\right)}{X^{2}}\,.$ (2.17)
Class II ($F\neq 0$, $A_{2}\neq A_{1}$)
In this class, $F\neq 0$ and $A_{2}\neq A_{1}$. There are two subclasses:
  • Class IIa [N-IIIi] ($F-XA_{1}\neq 0$))
    There are three free functions, $F$, $A_{1}$ and $A_{2}$. Also,
    $A_{3}$ $=\frac{4F_{X}(A_{1}+3A_{2})}{F}-\frac{2(A_{1}+4A_{2}-2F_{X})}{X}-\frac{4F}{X^{2}}\,,$ (2.18)
    $A_{5}$ $=\frac{2}{F^{2}X^{3}}\left[4F^{3}+F^{2}X(3A_{1}+8A_{2}-12F_{X})\right.$
    $\left.\qquad+8F\,F_{X}X^{2}(F_{X}-A_{1}-3A_{2})+6F_{X}^{2}X^{3}(A_{1}+3A_{2})\right]\,.$ (2.19)
  • Class IIb [N-IIIii] ($F-XA_{1}=0$))
    There are three free functions, $F$, $A_{2}$ and $A_{3}$ with $A_{4}$ and $A_{5}$ given by
    $A_{4}$ $=\frac{8F_{X}^{2}}{F}-\frac{4F_{X}}{X}\,,$ (2.20)
    $A_{5}$ $=\frac{1}{4FX^{3}(F+A_{2}X)}\left[FA_{3}^{2}X^{4}-4F^{3}-8F^{2}X(A_{2}-2F_{X})\right.$
    $\left.-4FX^{2}\left(4F_{X}\left(F_{X}-2A_{2}\right)+A_{3}F\right)+8F_{X}X^{3}(A_{3}F-4A_{2}F_{X})\right]\,.$ (2.21)
Class III ($F=0$)
This class corresponds to theories minimally coupled to gravity and can be split into three sub-classes.
  • Class IIIa [M-I] ($02A_{1}+A_{4}X=$ 0)
    There are three free functions $A_{1}$, $A_{2}$ and $A_{3}$, with $A_{5}$ given by
    $A_{5}=\frac{4A_{1}\left(A_{1}+2A_{2}\right)-4A_{1}A_{3}X+3A_{3}^{2}X^{2}}{4\left(A_{1}+3A_{2}\right)X^{2}}.$ (2.22)
  • Class IIIb [M-II] ($A_{1}+3A_{2}=0$)
    There are three free functions $A_{1}$, $A_{4}$ and $A_{5}$, with $A_{3}$ given by
    $A_{3}=\frac{2A_{1}}{3X}\,.$ (2.23)
  • Class IIIc [M-III] ($A_{1}=0$)
    There are four free functions, $A_{2}$, $A_{3}$, $A_{4}$ and $A_{5}$.
We note that the classes are not necessary disjoint.
3 Effective Field Theory of Dark Energy and cosmological constraints
In order to confront predictions of a large number of dark energy and modified gravity theories to observations, an effective approach has been developed: the Effective Field Theory (EFT) of dark energy [20, 21, 22, 23]. This approach, which extends the EFT of inflation [24, 25], is able to systematically treat all single-field models in the same framework, and hence to systematically constrain their parameters (see [26] for a review).
The metric is universally coupled to matter fields and the action is expressed in the most general unitary gauge consistent with the residual unbroken symmetries of spatial diffeomorphisms. The background evolution only depends on three operators, which all the other operators starting at least at quadratic order in the perturbations. Hence, their effects can be studied independently. The main formalism and results relevant for DHOST theories are presented in the next paragraphs.
The framework is based on the ADM 3+1 decomposition of spacetime [27], where spacetime is foliated into uniform scalar field hypersurfaces and expressed in terms of a lapse function $N$ and shift vector $N^{i}$,
$ds^{2}=-N^{2}dt^{2}+h_{ij}(dx^{i}+N^{i}dt)(dx^{j}+N^{j}dt)\,.$ (3.1)
Then the action can be expressed in terms of the extrinsic curvature tensor $K_{ij}$
$K_{ij}\equiv\frac{1}{2N}\left(\dot{h}_{ij}-D_{i}N_{j}-D_{j}N_{i}\right)\,,$ (3.2)
as
$S=\int d^{3}x\,dt\,N\sqrt{h}\,L[N,K_{ij},^{(3)}R_{ij};t]\,,$ (3.3)
where ${}^{(3)}R_{ij}$ is the Ricci tensor associated with the spatial metric $h_{ij}$.
To investigate the dynamics of linear perturbations, one needs to consider the perturbations of the quantities in the action
$\delta N\equiv N-\bar{N}\,,\qquad\delta K^{i}_{j}=K^{i}_{j}-H\delta^{i}_{j}\,,\qquad\delta^{(3)}R^{i}_{j}=^{(3)}R^{i}_{j}\,,$ (3.4)
where $H=\dot{a}/(\bar{N}a)$ is the Hubble parameter, and ${}^{(3)}R^{i}_{j}$ is a perturbation as ${}^{(3)}\bar{R}^{i}_{j}=0$. The Lagrangian at quadratic order in DHOST theories can be expressed as
$S_{\rm quad}$ $=$ $\int d^{3}x\,dt\,a^{3}\frac{M^{2}}{2}\bigg{\{}\delta K_{ij}\delta K^{ij}-\left(1+\frac{2}{3}\alpha_{L}\right)\delta K^{2}+(1+\alpha_{T})\bigg{(}\delta^{(3)}R\,\frac{\delta\sqrt{h}}{a^{3}}+\delta_{2}^{(3)}R\bigg{)}$ (3.7)
$\qquad\qquad\qquad\qquad+H^{2}\alpha_{K}\delta N^{2}+4H\alpha_{B}\delta K\delta N+({1+\alpha_{H}})\,^{(3)}R\,\delta N$
$\qquad\qquad\qquad\qquad+4\beta_{1}\,\delta K{\delta\dot{N}}+\beta_{2}\,{\delta\dot{N}}^{2}+\frac{\beta_{3}}{a^{2}}(\partial_{i}\delta N)^{2}\bigg{\}}\,,\qquad\quad$
where $\delta_{2}^{(3)}R$ represents the second order term in the perturbative expansion of ${}^{(3)}R$, and $M$, $\alpha_{L}$, $\alpha_{T}$, $\alpha_{K}$, $\alpha_{B}$, $\alpha_{H}$, $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$ are time-dependent functions.
Parameters $M$, $\alpha_{T}$, $\alpha_{K}$, $\alpha_{B}$ are specific to Horndeski theories, while the additional parameter $\alpha_{H}$ appears in beyond-Horndeski theories. The physical description of these parameters is discussed in [28].
Out of the four DHOST-specific parameters, only one is independent, and this is due to the degeneracy conditions. Hence, the DHOST theories can be classified as:
$\mathcal{C}_{\rm I}$ $:\quad\alpha_{L}=0\,,\qquad\beta_{2}=-6\beta_{1}^{2}\,,\qquad\beta_{3}=-2\beta_{1}\left[2(1+\alpha_{H})+\beta_{1}(1+\alpha_{T})\right]\,,$ (3.8)
$\mathcal{C}_{\rm II}$ $:\quad\beta_{1}=-(1+\alpha_{L})\frac{1+\alpha_{H}}{1+\alpha_{T}}\,,\quad\beta_{2}=-6(1+\alpha_{L})\frac{(1+\alpha_{H})^{2}}{(1+\alpha_{T})^{2}}\,,\quad\beta_{3}=2\frac{(1+\alpha_{H})^{2}}{1+\alpha_{T}}\,.$ (3.9)
In order to investigate the stability of the solutions, one can derive the quadratic tensor action
[…]
Hence, the parameter $\alpha_{T}$ is associated to the tensor propagating velocity,
$c_{T}^{2}=1+\alpha_{T}\,.$ (3.11)
For the scalar, that is
[…]
For $\mathcal{C}_{\rm I}$ theories, $A_{\tilde{\zeta}}$ and ${B}_{\tilde{\zeta}}$ are given by
${A}_{\tilde{\zeta}}$ $=\frac{1}{(1+\alpha_{B}-\dot{\beta}_{1}/H)^{2}}\bigg{[}\alpha_{K}+6\alpha_{B}^{2}-\frac{6}{a^{3}H^{2}M^{2}}\frac{d}{dt}\left(a^{3}HM^{2}\alpha_{B}\beta_{1}\right)\bigg{]}\;,$ (3.13)
${B}_{\tilde{\zeta}}$ $=-2(1+\alpha_{T})+\frac{2}{aM^{2}}\frac{d}{dt}\bigg{[}\frac{aM^{2}\big{(}1+\alpha_{H}+\beta_{1}(1+\alpha_{T})\big{)}}{H(1+\alpha_{B})-\dot{\beta}_{1}}\bigg{]}\;.$ (3.14)
The stability conditions require
${A}_{\tilde{\zeta}}>0\,,\qquad{B}_{\tilde{\zeta}}>0\,.$ (3.15)
In the case of $\mathcal{C}_{\rm II}$ theories ${B}_{\tilde{\zeta}}=-2(1+\alpha_{T})$, which combined with the result for the tensor modes, shows that scalar and tensor modes cannot be stable simultaneously, and hence this class of theories is not viable.
The effects of the gravitational wave event GW170817 [29], observed by the LIGO/Virgo collaboration, and of the gamma-ray burst GRB 170817A [30] on the DHOST models have been discussed in [31, 32]. Fixing the gravitational wave speed to be equal to the speed of light only allows a Lagrangian of the form
$L^{{}_{\rm DHOST}}_{c_{g}=1}=P+Q\Box\phi+F\,{}^{(4)}\!R+A_{3}\phi^{\mu}\phi^{\nu}\phi_{\mu\nu}\Box\phi$
$+\frac{1}{8F}\bigg{(}48F_{X}^{2}-8(F-XF_{X})A_{3}-X^{2}A_{3}^{2}\bigg{)}\phi^{\mu}\phi_{\mu\nu}\phi_{\lambda}\phi^{\lambda\nu}$
$+\frac{1}{2F}\left(4F_{X}+XA_{3}\right)A_{3}(\phi_{\mu}\phi^{\mu\nu}\phi_{\nu})^{2}\;.$ (3.16)
This restriction is based on the assumption that scales probed by the LIGO-Virgo measurements (which are of the order of $10^{3}$ km), can be extrapolated by several orders of magnitude to cosmological scales.
The cosmological evolution of these models has been discussed in [33, 34, 35, 36, 37]. Some of the studies include a relaxation of the assumption related to the tensor mode speed, allowing much more general DHOST theories.
The EFT method has been used, together with the CMB likelihoods from Planck, to place constraints on the EFT parameters in DHOST models [38, 39].
4 Astrophysical bodies and the Vainshtein mechanism
Significant work has been done on the effect of DHOST theories on astrophysical objects, in particular gravitational screening, neutron stars and black holes.
4.1 The Vainshtein mechanism
Dark energy and modified gravity models can change the spacetime evolution compared to general relativity, but these modifications should not appear at small scales, where precise laboratory and solar system tests can be performed. This can be achieved by the so-called screening mechanisms. In the case of DHOST theories, the Vainshtein screening [40] is at play. It has been shown to work well for Horndeski theories [41, 42, 43], whereas in the case of beyond-Horndeski theories the standard gravity is modified inside matter [44]. This new behaviour has been studied for various astrophysical objects, as presented in [45]. For DHOST theories, this topic is discussed in general in [31, 46, 47], with specific examples in [48, 49, 50, 51].
To illustrate how the Vainshtein screening mechanism works in the context of DHOST theories, one considers a non-relativistic spherical object of density $\rho(r)$, which induces a perturbation of the metric, expressed in terms of the Newtonian potentials $\Phi$ and $\Psi$ as
$ds^{2}=-\left[1+2\Phi(r)\right]dt^{2}+\left[1-2\Psi(r)\right]\delta_{ij}\,dx^{i}dx^{j}\,.$ (4.1)
The spherical object also perturbs the scalar field,
$\phi=\phi_{c}(t)+\chi(r)\,,$ (4.2)
and the mass inside the radius $r$ is
[…]
The two Newtonian potentials satisfy the differential equations [47]
$\Phi^{\prime}(r)$ $=G_{\rm N}\left(\frac{m(r)}{r^{2}}+\Xi_{1}{m^{\prime\prime}(r)}\right)\,,$ (4.4)
$\Psi^{\prime}(r)$ $=G_{\rm N}\left(\frac{m(r)}{r^{2}}+\Xi_{2}\frac{m^{\prime}(r)}{r}+\Xi_{3}{m^{\prime\prime}(r)}\right)\,,$ (4.5)
where
$\Xi_{1}$ $\equiv\frac{(\alpha_{H}+c_{T}^{2}\beta_{1})^{2}}{c_{T}^{2}(1+\alpha_{V}-4\beta_{1})-\alpha_{H}-1}\,,$ (4.6)
$\Xi_{2}$ $\equiv-\frac{\alpha_{H}(\alpha_{H}-\alpha_{V}+2(1+c_{T}^{2})\beta_{1})+\beta_{1}(c_{T}^{2}-1)(1+c_{T}^{2}\beta_{1})}{c_{T}^{2}(1+\alpha_{V}-4\beta_{1})-\alpha_{H}-1}\,,$ (4.7)
$\Xi_{3}$ $\equiv-\frac{\beta_{1}(\alpha_{H}+c_{T}^{2}\beta_{1})}{c_{T}^{2}(1+\alpha_{V}-4\beta_{1})-\alpha_{H}-1}\,.$ (4.8)
These coefficients quantify the breaking of the Vainshtein screening inside matter and depend on the DHOST parameters.
Outside the source, $m^{\prime}=m^{\prime\prime}=0$ and hence the standard Newtonian potential is recovered $\Phi=\Psi=G_{\rm N}m/r$. Inside the source, the behaviour is modified, similarly to the beyond-Horndeski scenario [44].
The screening in DHOST theories has been studied in the context of the LIGO/Virgo observations in [52], and stringent constraints on the $\beta_{1}$ parameter are placed. In [48] a mechanism on evading these gravitational wave constraints is proposed.
4.2 Newtonian and Neutron stars
These results can be used to investigate the impact of such theories on various astrophysical bodies. In the case of Newtonian stars, there is a theoretical upper bound for function $\Xi_{1}$ [53].
$\Xi_{i}>-\frac{1}{6}$ (4.9)
This quantity can also be constrained from observations, including the lowest mass hydrogen burning stars [54], white dwarfs [55] and by analysing weak lensing and X-ray profiles of galaxy clusters [56, 57].
Neutron stars have been studied in a subclass of DHOST models, depending on a single free function $F(X)$ [58],
$S_{\rm grav}=\int d^{4}x\,\sqrt{-g}\left(F\,R-4\frac{F_{X}}{X}(L_{3}-L_{4})+L_{\rm m}\right)\,.$ (4.10)
For non-rotating neutron stars, the metric is given by
$ds^{2}=-f(r)dt^{2}+h(r)dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\!\theta\,d{\phi}^{2}\right).$ (4.11)
Assuming a spherically symmetric scalar field solution of the form [59]
$\varphi(t,r)=q\;t+\psi(r)\,,$ (4.12)
the action (4.11) becomes
$S=\int d^{4}x\,r^{2}\sqrt{fh}\left(\frac{2Fh^{\prime}}{h^{2}r}+\frac{2F(h-1)}{hr^{2}}-\frac{4F_{X}X^{\prime}\left(fX+q^{2}\right)}{fhrX}+L_{m}\right)\,,$ (4.13)
where
$X=\psi^{\prime 2}/h-q^{2}/f\,.$ (4.14)
Here the matter appearing the Lagrangian $L_{m}$ is modelled by a perfect fluid with energy density $\rho$, pressure $P$ and 4-velocity $u^{\mu}$. The energy-momentum tensor is thus
$T^{\mu\nu}=\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}\,L_{m})}{\delta g_{\mu\nu}}=(\rho+P)u^{\mu}u^{\nu}+Pg^{\mu\nu}\,.$ (4.15)
From the energy-momentum tensor one obtains
$P^{\prime}=-\frac{f^{\prime}(P+\rho)}{2f}\,.$ (4.16)
Considering a Taylor expansion of the function $F$ of the form,
$F=\frac{\kappa}{2}+\sigma X\,,$ (4.17)
the action (4.13) can be varied to find solutions for $\psi$. Then redefining $\psi$ as
$\psi\to\frac{p^{1/2}}{q}\psi\,,\qquad p\equiv q^{2}\sigma\,,$ (4.18)
one gets
$\psi^{\prime 2}=\Delta_{1}-\Delta_{2},$ (4.19)
where
$\Delta_{1}$ $=$ $\frac{h\left[f\left(2\kappa+r^{2}P\right)+8p\right]-4p}{8f}\,,$ (4.20)
$\Delta_{2}^{2}$ $=$ $\Delta_{1}^{2}-\frac{hp\left[f\left(h\left(2\kappa+2Pr^{2}+\rho r^{2}\right)-2\kappa\right)+4p\left(h-1\right)\right]}{4f^{2}}\,.$ (4.21)
The deviations from general relativity $\Xi_{1}$ and $\Xi_{2}$ are given by [31]
$\Xi_{1}=-2\,\Xi_{2}=\frac{2p}{2p+\kappa}.$ (4.22)
Outside the star $X=-q^{2}$ and hence the modification is indistinguishable from general relativity. Inside the star, their radial profiles must be considered, which depend on their central density and on the equation of state. This can be achieved by numerically integrating the equations of motion. In [58] several scenarios have been considered and it is shown that in DHOST theories neutron stars can have masses and radii significantly larger than in general relativity, which could provide explanations for several observations, such as the pulsar PSR J1614-2230 [60] or the compact object measured by GW190814 event [61], which have these properties.
These results were extended to slowly rotating stars [62, 63]. The metric (4.10) is generalised to
$ds^{2}=-f(r)dt^{2}+h(r)dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\!\theta\,d{\phi}^{2}\right)+2w(r,\theta)r^{2}\sin^{2}\!\theta\,dtd\phi\,.$ (4.23)
The function $w$, related to rotation, is assumed to be small, i.e. $w(r,\theta)\ll f$. The perfect fluid in the star is also rotating, which can be described by the four-velocity
[…]
up to first order in the fluid angular velocity $\Omega=d\phi/dt=u^{\phi}/u^{t}$. The second order action analogous to (4.13) is
$S^{(2)}=\int d^{4}x\sin^{3}\theta\left(\frac{Fr^{4}}{2\sqrt{f}\sqrt{h}}(\partial_{r}w)^{2}+\frac{F\sqrt{h}r^{2}}{2\sqrt{f}}(\partial_{\theta}w)^{2}+\frac{\sqrt{h}\rho r^{4}}{2\sqrt{f}}w^{2}+L_{m}^{(2)}\right)\,,$ (4.25)
which yields the equations of motion
$\partial_{rr}w+\left(\frac{4}{r}-\frac{f^{\prime}}{2f}-\frac{h^{\prime}}{2h}+\frac{F_{X}X^{\prime}}{F}\right)\partial_{r}w+\frac{h}{r^{2}}\left(\partial_{\theta\theta}w+3\cot\theta\,\partial_{\theta}w\right)$
$-\frac{h(P+\rho)}{F}(\Omega+w)=0\,.$ (4.26)
These equations can be expressed as an infinite sum of Legendre polynomials
$w(r,\theta)=\sum_{l}w_{l}(r)\,P_{l}(\theta)\,.$ (4.27)
Outside the star, the results are the same as in general relativity. Inside the star, the equation can be solved numerically for each $l$ and related to the angular momentum.
4.3 Black Holes
A renewed interest in black holes can be attributed to the LIGO/Virgo direct detection of gravitational waves from binary black hole and neutron star mergers. Black holes can also be used as probes to investigate deviations from general relativity. Hence, one has to make sure that theories of modified gravity or dark energy yield constraints compatible with black holes. In standard general relativity, the most general axisymmetric stationary rotating black hole solution is the Kerr solution [64, 65]. Indeed, in the binary black hole observation, it is thought that the merger involved two spinning Kerr black holes [66]. These Kerr black holes can be generalised for DHOST theories and solutions that have non-trivial scalar field (stealth hair) have been derived [67]; linear perturbations are discussed in [68]. Generating methods for solutions by using disformal transformations are discussed in [69, 70], or by using perturbations about the flat Minkowski metric $\eta_{\mu\nu}$ along a null direction (the Kerr-Schild ansatz) in [71, 72].
We present here the Kerr solutions from DHOST theories, and we mention at the end of the section other recent developments in the field. Quadratic shift-symmetric DHOST theories with $c_{T}=1$, where all the functions in the Lagrangian only depend on $X$ are considered, such that $X=X_{0}$ is a constant. Moreover, stealth solutions are sought (solutions where the metric $g_{\mu\nu}$ is also a solution of the vacuum Einstein equations with a cosmological constant). Although such solutions are plagued by the strong coupling problem [73, 74], a controlled detuning, the scordatura mechanism [75], can be employed to solve the issue. The assumptions imply that
$P+2\Lambda F=0\,,\quad P_{X}+\Lambda(4F_{X}-X_{0}A_{1X})=0\,,\quad Q_{X}=0\,,\quad A_{1}=0\,\quad A_{3}+2A_{1X}=0\,,$ (4.28)
where the functions are evaluated at $X=X_{0}$.
By considering solutions that satisfy $c_{T}=1$, one obtains
$A_{1}=A_{2}=0\,,\qquad Q=0\,,$
$P(X_{0})+2\Lambda F(X_{0})=0\,,\quad P_{X}(X_{0})+4\Lambda F_{X}(X_{0})=0\,,\quad A_{3}(X_{0})=0\,.$ (4.29)
The metric that is obtained is the usual Kerr solution in general relativity, which can be expressed in Boyer-Lindquist coordinates $(t,r,\theta,\psi)$, as
$ds^{2}=-\frac{\Delta_{r}}{\Xi^{2}\rho}\left(dt-a\sin^{2}{\theta}\,d\psi\right)^{2}+\rho\left(\frac{dr^{2}}{\Delta_{r}}+\frac{d\theta^{2}}{\Delta_{\theta}}\right)+\frac{\Delta_{\theta}\sin^{2}{\theta}}{\Xi^{2}\rho}\left(adt-\left(r^{2}+a^{2}\right)d\psi\right)^{2}\,,$ (4.30)
where the cosmological constant $\Lambda=3/\ell^{2}$, $\Xi\equiv 1+{a^{2}}/{\ell^{2}}$ is a constant, and
$\Delta_{r}$ $=\left(1-\frac{r^{2}}{\ell^{2}}\right)\left(r^{2}+a^{2}\right)-2Mr\;,\qquad\Delta_{\theta}=1+\frac{a^{2}}{\ell^{2}}\cos^{2}{\theta}\;,\quad\rho=r^{2}+a^{2}\cos^{2}{\theta}\,.$ (4.31)
$M$ is the mass of the black hole and $a$ the angular momentum parameter satisfying the condition $a\leq M$.
These stealth black holes also have scalar hair characterised by [67]
[…]
where $E$ is a constant and
$\mathcal{R}(r)\equiv m^{2}\left(r^{2}+a^{2}\right)\left[\eta^{2}\left(r^{2}+a^{2}\right)-\Delta_{r}\right]\;,\qquad\Theta(\theta)\equiv a^{2}m^{2}\sin^{2}{\theta}\left(\Delta_{\theta}-\eta^{2}\right)\,,$ (4.33)
with $X_{0}=-m^{2}$ and $\eta\equiv\Xi E/m$. If there is no cosmological constant ($\ell\to\pm\infty$), then $\eta=1$, $\Theta=0$, $E=m$ and
$\mathcal{R}(r)\equiv 2Mm^{2}r\left(r^{2}+a^{2}\right)\;.$ (4.34)
New Kerr solutions can be derived by considering disformal transformations of the metric of the form
$\tilde{g}_{\mu\nu}=A(X)g_{\mu\nu}+B(X)\phi_{\mu}\phi_{\nu}\,,$ (4.35)
where $A$ and $B$ are the disformal potentials, depending on $X$. DHOST theories are stable under these transformations [13]. Hence, assuming that the tilded action satisfies the stealth conditions (4.28), with a Kerr solution (4.32), the conditions for the untilded conditions become
$P+\frac{2\Lambda N}{A_{0}}F=0\,,\qquad{\partial_{X}}\left[P+\frac{\Lambda}{A_{0}}\left(4+\frac{B_{0}X_{0}}{N}\right)F-\frac{\Lambda X_{0}}{N^{3}}A_{1}\right]=0\,,$ (4.36)
$Q_{X}=0\,,\qquad A_{1}-\frac{N^{2}B_{0}}{A_{0}}F=0\,,$ (4.37)
$\frac{A_{0}}{2N}A_{3}+\left(2B_{0}N_{X}+\frac{B_{0}N^{2}}{A_{0}}\right)F+2B_{0}NF_{X}+B_{0}NA_{2}+\frac{N^{8}}{A_{0}}\partial_{X}\left(\frac{A_{1}}{N^{3}}-\frac{B_{0}}{A_{0}}\frac{F}{N}\right)=0\,,$ (4.38)
where the disformal transformation is constant ($A(X)=A(X_{0})\equiv A_{0}$, $B(X)=B(X_{0})\equiv B_{0}$) and the derivatives are evaluated at $X_{0}$. Imposing (4.3), these become
$A_{1}=-A_{2}=N^{2}\frac{B_{0}}{A_{0}}F\,,\qquad Q=0\,,$ (4.39)
$\partial_{X}\left(P+\frac{4\Lambda}{A_{0}}F\right)=0\,,\qquad\frac{A_{0}}{2NB_{0}}A_{3}+\left(2N_{X}+\frac{N^{2}}{A_{0}}(1-NB_{0})\right)F+2NF_{X}=0\,.$ (4.40)
In this case, the untilded theory admits the disformed Kerr black hole solution. Fixing $A_{0}=1$, the metric becomes
$g_{\mu\nu}\;=\;\tilde{g}_{\mu\nu}-B_{0}\,\phi_{\mu}\phi_{\nu}\,,$ (4.41)
which leads to a new Kerr solution. To avoid pathological behaviour at $r\to\infty$ (non-asymptotically flat metric at infinity), a vanishing cosmological constant is required, implying $\eta=1$, $\Theta=0$ and $E=m$ as before.
The scalar field solution becomes
[…]
and the new solution is
$ds^{2}$ $=-\frac{\Delta_{r}}{\rho}\left(dt-a\sin^{2}{\theta}\,d\psi\right)^{2}+\frac{\rho}{\Delta_{r}}{dr^{2}}+\rho\,{d\theta^{2}}+\frac{\sin^{2}{\theta}}{\rho}\left(a\,dt-\left(r^{2}+a^{2}\right)d\psi\right)^{2}$
$+\alpha\left(dt\pm{\sqrt{2Mr(r^{2}+a^{2})}}/{\Delta_{r}}\,dr\right)^{2}\,,$ (4.43)
with $\alpha\equiv-B_{0}m^{2}$. The inverse of this metric is
$g^{\mu\nu}\,=\,\tilde{g}^{\mu\nu}+\frac{\alpha}{m^{2}(1-\alpha)}\phi^{\mu}\phi^{\nu}\,,$ (4.44)
where $\tilde{g}^{\mu\nu}$ is the inverse Kerr metric while the only non-vanishing components of $\phi^{\mu}=\tilde{g}^{\mu\nu}\phi_{\nu}$ are
$\phi^{t}=\frac{m}{\Delta_{r}}\left(r^{2}+a^{2}+\frac{2Mra^{2}\sin^{2}\theta}{\rho}\right)\,,\quad\phi^{r}=m\frac{\sqrt{2Mr(r^{2}+a^{2})}}{\rho}\,,\quad\phi^{\varphi}=m\frac{2aMr}{\Delta_{r}\rho}\,.$ (4.45)
Hence, the disformal transformation of the stealth Kerr solution considered yields a non-stealth exact solution which is parametrised by an additional deformation parameter $\alpha$ which encodes precisely the deviations from general relativity, in addition to the mass and angular momentum parameters $(M,a)$ of the Kerr family. The sign ambiguity in (4.3) can be absorbed by redefining $t\to\pm t$ and $a\to\pm a$.
At infinity where $r\to+\infty$, the disformed Kerr metric becomes equivalent to
$ds^{2}$ $\simeq-\left(1-\frac{2M_{1}}{r}\right)dt^{2}+\left(1-\frac{2M_{2}}{r}\right)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta\,d\varphi^{2})$
$+2\alpha\sqrt{\frac{2M_{1}}{r}}dr\,dt+{\cal O}\left(\frac{1}{r^{2}}\right)\,,$ (4.46)
where
$M_{1}\equiv\frac{M}{1-\alpha}\,,\quad M_{2}\equiv(1+\alpha)M\,.$ (4.47)
This new solution is not circular [76], but similar to other Kerr black holes, it does not have a physical singularity.
In [77] it is shown that LISA can be in principle used to distinguish the Kerr geometry in DHOST from general relativity by measuring the deformation parameter $\alpha$ which quantifies departures from the standard Kerr black hole spacetime through an accumulated effect. This parameter can also be investigated by looking at its effect on the black hole shadow [78] using the observations from the Event Horizon Telescope [79, 80].
A different class of black holes solutions is the non-rotating Schwarzschild one, which can be obtained from quadratic DHOST theories [81, 74, 82], but some are disfavoured through a perturbation analysis. Perturbations to such black holes are investigated in [83, 84, 85, 86]. The strong coupling of the solutions in the DHOST theories is discussed [75], together with a solution – adding small corrections (scordatura) to avoid this issue and the implication for the black hole solutions [87]. Various generalisations have been considered, e.g. non-Schwarzschild solutions with a non-constant kinetic term [88].
Black hole solutions have been studied [89] for shift-symmetric DHOST theories which admit exact black hole solutions with a linearly time-dependent scalar field. In this case, only purely quadratic theories are viable and exact static and spherically symmetric (Schwarzschild and Schwarzschild-(anti-)de Sitter) are determined.
In [90, 91] the no-hair theorem [92] is generalised to DHOST theories, showing that the coupling with Gauss-Bonnet is necessary to have hair.
4.4 Other astrophysical constraints
In this section we briefly introduce some of the additional effects of DHOST theories from the late universe that have been studied and that can lead to observable effects
  • The modification of the sound speed in the atmosphere of the Earth, due to the significant modification of the propagation of matter perturbations.[93]. This was used to constrain the parameters of the DHOST model.
  • Constraints on DHOST models from weak lensing, X-ray and Sunyaev - Zel’dovich observables [94, 95].
  • Constraints from large-scale structure, using two and three-point statistics with the BOSS data [96].
  • The use of helioseismology to place constraints on the parameters of the DHOST models, using the Sun as a laboratory [97, 98].
  • The effect of DHOST on structure formation through modifications in the matter power spectrum [99] and the increased sensitivity on the short-wavelength behaviour of the linear power spectrum.
5 Inflation
The inflationary paradigm, initially designed solve several cosmological paradoxes, including the horizon problem, the flatness problem and the magnetic monopole problem, is also responsible for the formation of the structures through the growth of primordial quantum fluctuations. Cosmic Microwave background (CMB) experiments, such as COBE [100], WMAP [101] and Planck [102] have tightened constraints on the observable effects on inflation, for which the main constraints are based on measurements of the angular power spectrum of CMB anisotropies and on adiabatic scalar perturbations. This exhibits a small departure from scale invariance that can be quantified at the power spectrum level through
$\mathcal{P}_{\zeta}(k)=\mathcal{P}_{\zeta}(k_{*})\left(\frac{k}{k_{*}}\right)^{n_{s}-1}\,,$ (5.1)
where $k_{*}$ is a pivot scale and $n_{s}$ is the scalar spectral index. Further constraints can be determined by expanding this further as
$\log\mathcal{P}_{\zeta}(k)=\log\mathcal{P}_{\zeta}(k_{*})+\frac{1}{2}\frac{d\log n_{s}}{d\log k}\left(\frac{k}{k_{*}}\right)^{2}+\frac{1}{6}\frac{d^{2}\log n_{s}}{d\log k^{2}}\left(\frac{k}{k_{*}}\right)^{3}+\ldots\,.$ (5.2)
Planck has placed stringent constrains on these coefficients,
$n_{s}$ $=0.9625\pm 0.0048\,,$ (5.3)
$\alpha_{s}$ $=\frac{dn_{s}}{d\ln k}=0.002\pm 0.010\,,$ (5.4)
$\beta_{s}$ $=\frac{d\alpha_{s}}{d\ln k}=0.010\pm 0.013\,,$ (5.5)
$\ln(10^{10}A_{s})$ $=3.044\pm 0.014\,$ (5.6)
at $68\%$ confidence level, using the TT,TE,EE+lowE+lensing likelihoods [102] at a pivot scale of $k_{*}=0.05\mathrm{~{}Mpc}^{-1}$.
An additional constraint comes from the tensor-to-scalar ratio, currently bounded at the level of $r<0.044$ at 95% confidence [103] when combining Planck with BICEP2/Keck 2015 data [104].
Further constraints can be determined by investigating predictions on higher-order correlators, such as the bispectrum (three-point correlation function). In particular, Planck has placed the tightest constraints so far on the amplitudes of non-Gaussianities, which are three-dimensional quantities. Based on the model of inflation considered, they come predominantly in the following three shapes: local, equilateral and orthogonal, which can be expressed in terms of the primordial power spectrum $P_{\Phi}$ as
$B_{\Phi}^{\text{loc}}$ $(k_{1},k_{2},k_{3})=2\left[P_{\Phi}(k_{1})P_{\Phi}(k_{2})+\text{2perms}\right]\,,$ (5.7)
$B_{\Phi}^{\text{equil}}$ $(k_{1},k_{2},k_{3})=6\left\{-[P_{\Phi}(k_{1})P_{\Phi}(k_{2})+\text{2 perms}]\right.$
$-2[P_{\Phi}(k_{1})P_{\Phi}(k_{2})P_{\Phi}(k_{3})]^{2/3}+[P_{\Phi}^{1/3}(k_{1})P_{\Phi}^{2/3}(k_{2})P_{\Phi}(k_{3})+\text{5 perms}]\left.\right\}\,,$ (5.8)
$B_{\Phi}^{\text{orth}}$ $(k_{1},k_{2},k_{3})=6\big{[}3(P_{\Phi}^{1/3}(k_{1})P_{\Phi}^{2/3}(k_{2})P_{\Phi}(k_{3})+5\text{ perms})$
$-3\left[P_{\Phi}(k_{1})P_{\Phi}(k_{2})+\text{2 perms}\right]-8(P_{\Phi}(k_{1})P_{\Phi}(k_{2})P_{\Phi}(k_{3}))^{2/3}\big{]}\,.$ (5.9)
For a bispectrum of one of these shapes, $f_{\rm NL}$ represents the amplitude of its non-Gaussianities. The current Planck constraints on these shapes are
$-11.1<f_{\rm NL}^{\rm local}$ $<9.3$ (5.10)
$-120<f_{\rm NL}^{\rm equil}$ $<68$ (5.11)
$-86<f_{\rm NL}^{\rm orth}$ $<10$ (5.12)
at 95 %CL [105].
As mentioned in the introduction, theories designed to explain dark energy can also be adequate to describe inflation. In the case of DHOST theories, work has been performed to build models compatible with the constraints on inflation from both power spectrum [106] and bispectrum [107, 108, 109].
The starting point is a simplification of the quadratic DHOST action,
$S_{\rm D}=\int d^{4}x\sqrt{-g}\bigg{[}F_{0}(X)+F_{1}(X)\Box\phi+F_{2}(X)R+\frac{6F_{2,X}^{2}}{F_{2}}\phi^{\nu}\phi_{\nu\eta}\phi^{\eta\lambda}\phi_{\lambda}\bigg{]}\,,$ (5.13)
where an assumption has been made that the functions the $F_{i}$ and $A_{i}$ only depend on the kinetic term $X$.
This action can be used to determine the power spectrum of primordial curvature perturbations in the DHOST theories, using the methods developed in [75].
It turns out that inflationary solutions in a de Sitter phase of the universe would lead either to infinite strong couplings or gradient instability for scalar field perturbations. This problem can be rectified by slightly breaking the degeneracy conditions, adding scordatura corrections of the form $\alpha L_{i}^{(2)}$, for one $i=1,\dots,5$ (see Sec. 4.3). In this case, the theory is shift-symmetric (invariant under the symmetry $\phi\to\phi+c$). It can be shown that such theories lead to scale-invariant power spectra, that are not compatible with observations ($n_{s}=1$ is not allowed by CMB data).
To resolve this issue, one can consider Eq. (5.13) as the background and an axion-like potential [110] as a perturbation
$S_{\rm V}=-\int d^{4}x\sqrt{-g}\mu^{4}\left(\cos\frac{\phi}{f}-1\right)\approx\int d^{4}x\sqrt{-g}\bigg{[}-\frac{m_{\rm phys}^{2}}{2}\phi^{2}-\frac{\lambda_{\rm phys}}{4!}\phi^{4}\bigg{]}\,,$ (5.14)
where
$m^{2}_{\rm phys}=-\frac{\mu^{4}}{f^{2}},\ \lambda_{\rm phys}=\frac{\mu^{4}}{f^{4}}\,,$ (5.15)
for $\phi\ll f$.
The determination of the power spectrum has been performed in [106], based on [75], following the standard formalism of inflationary perturbation theory [111]. Here we illustrate the procedure for the Lagrangian without the perturbations $m^{2}$ and $\lambda$, with the full details given in [106]. The action is perturbed at first order in perturbations and expanded up to quadratic order, and then after suitable integrations by parts, the second order DHOST Lagrangian is obtained in terms of the comoving curvature perturbation $\zeta$,
~${\tilde{\cal L}}_{\rm D}^{(2)}=a^{3}f_{2}\bigg{(}\bar{\cal A}\,\dot{\zeta}^{2}-\bar{\cal B}\,\frac{{\tilde{k}}^{2}}{a^{2}}\zeta^{2}\bigg{)}\,,$ (5.16)
with
$\bar{\cal A}=6\bigg{[}1-\frac{\beta_{K}}{(1+\alpha_{B})^{2}}\bigg{]}\,,\hskip 28.45274pt\bar{\cal B}=-2\bigg{[}1-\frac{1}{af_{2}}\frac{d}{d{\tilde{t}}}\bigg{(}\frac{af_{2}}{h_{b}}\frac{1+\alpha_{H}}{1+\alpha_{B}}\bigg{)}\bigg{]}\,.$ (5.17)
Here the parameters $\alpha_{i}$ and $\beta_{i}$ depend on the functions $f_{i}$ and their derivatives and are given by
$\alpha_{H}\equiv-{\rm x}\frac{f_{2,{\rm x}}}{f_{2}}\,,\hskip 28.45274pt\alpha_{B}\equiv\frac{1}{2}\frac{\dot{\varphi}\,{\rm x}}{h_{b}}\frac{f_{1,{\rm x}}}{f_{2}}+\alpha_{H}\,,\hskip 28.45274pt\alpha_{K}\equiv-\frac{{\rm x}}{6h_{b}^{2}}\frac{f_{0,{\rm x}}}{f_{2}}+\alpha_{H}+\alpha_{B}\,,$ (5.18)
$\beta_{K}\equiv-\frac{{\rm x}^{2}}{3}\frac{f_{0,{\rm x}{\rm x}}}{h_{b}^{2}f_{2}}+(1-\alpha_{H})(1+3\alpha_{B})+\beta_{B}+\frac{(1+6\alpha_{H}-3\alpha_{H}^{2})\alpha_{K}-2(2-6\alpha_{H}+3\alpha_{K})\beta_{H}}{1-3\alpha_{H}}\,,$
$\beta_{B}\equiv\dot{\varphi}\,{\rm x}^{2}\frac{f_{1,{\rm x}{\rm x}}}{h_{b}f_{2}}\,,\hskip 28.45274pt\beta_{H}\equiv{\rm x}^{2}\frac{f_{2,{\rm x}{\rm x}}}{f_{2}}\,.$ (5.19)
This second-order action can be used to derive the power spectrum through field quantisation [112, 113, 114]. To examine the creation of primordial fluctuations from the Bunch-Davies vacuum in these models, it is convenient to parameterise the de Sitter phase in the early universe. Hence, the Lagrangian (5.16) can be used to express the second-order action as
[…]
where primes denote derivative with respect to conformal time; the Mukhanov-Sasaki variable $v_{k}$, defined as $v=z\zeta$, satisfies
$v_{k}^{\prime\prime}+\left(k^{2}c_{s}^{2}-\frac{z^{\prime\prime}}{z}\right)v_{k}=0\,,$ (5.21)
This equation can be solved using appropriate conditions at $\eta\to-\infty$ and the power spectrum of (scalar) curvature perturbations is computed as
$P_{\zeta}(k)=\frac{k^{3}}{2\pi^{2}}\frac{|v_{k}(\eta_{E})|^{2}}{z(\eta_{E})^{2}}\,,$ (5.22)
where $\eta_{E}$ is the time corresponding to the horizon entry.
During the inflationary epoch, as solutions linear in time are sought, $X=0$ and hence the $\alpha_{i}$ and $\beta_{i}$ functions become constant. There are five constant functions (and an additional two perturbation parameters) that can be chosen to match the CMB observations.
In the case of the bispectrum [107], one can employ the formalism described in [115, 116]: expanding the action at third order in the comoving gauge, computing the Hamiltonian in the interaction picture and employing the in-in formalism to determine the three-point correlation function. In addition to the parameters fit for the power spectrum, there are another five that appear in the bispectrum.
For the power spectrum, parameters that match the CMB observations can be readily obtained, but for the bispectrum this is more involved due to its three-dimensional nature, requiring a full CMB likelihood analysis to confront to the Planck data. Parameters compatible with both the powers spectrum and the bispectrum have been found in [108].
6 Alternatives to inflation
Another idea related to the use of DHOST models in the early universe is presented in [117, 118, 119, 120], where an alternative to inflation – the bouncing cosmology [121, 122, 123, 124] – is built using these theories.
This model is inspired by the Galilean Genesis model [125, 126] and aims to resolve several potential problems such as no-go theorems, strong coupling issues and connections to standard cosmology.
In this model, the starting action is
$S=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+\mathcal{L}_{H2}+\mathcal{L}_{H3}+\mathcal{L}_{D}\right)~{},$ (6.1)
where $\mathcal{L}_{H2}$ and $\mathcal{L}_{H3}$ and $\mathcal{L}_{D}$ are given by
$\mathcal{L}_{H2}$ $=K(\pi,X)=-g_{1}(\pi)X+g_{2}(\pi)X^{2}~{},$ (6.2)
$\mathcal{L}_{H3}$ $=G(X)\Box X~{},~{}G(X)=\gamma X^{2}~{},$ (6.3)
$\mathcal{L}_{D}$ $=\frac{R}{2}h-\frac{h}{4X}\left[\pi_{\mu\nu}\pi^{\mu\nu}-(\Box\pi)^{2}\right]+\frac{h-2Xh_{X}}{4X^{2}}\left[\pi_{\mu}\pi^{\mu\rho}\pi_{\rho\nu}\pi^{\nu}-(\Box\pi)\pi^{\mu}\pi_{\mu\nu}\pi^{\nu}\right]\,,$ (6.4)
where
$g_{1}(\pi)=\frac{3}{2}f^{2}e^{4\pi}\frac{1+e^{2\pi}}{1+e^{m\pi}}~{},~{}g_{2}(\pi)=e^{2\pi}\frac{1+e^{2\pi}}{1+e^{4\pi}}~{},$ (6.5)
$h(X)=d_{1}X+d_{2}X^{2}\,.$ (6.6)
The parameter $m>6$ determines the behaviour when the universe exiting the quasi-Minkowskian state and $m=7$ through is used in the analysis.
For the background, the Friedmann equations are obtained
$3H^{2}=\rho_{\pi}~{},~{}-2\dot{H}=\rho_{\pi}+p_{\pi}~{},$ (6.7)
where $H$ is the Hubble parameter, $\rho_{\pi}$ and $p_{\pi}$ are the density and pressure of the matter field $\pi$, expressed by
$\rho_{\pi}=-\frac{1}{2}g_{1}\dot{\pi}^{2}+\frac{3}{4}g_{2}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},$ (6.8)
$p_{\pi}=-\frac{1}{2}g_{1}\dot{\pi}^{2}+\frac{1}{4}g_{2}\dot{\pi}^{4}-\gamma\dot{\pi}^{4}\ddot{\pi}\,,$ (6.9)
independent of the DHOST term (6.4) [117, 118]. Assuming as the initial condition of scalar field $\pi$ at far past to be $\pi\ll-1$, the first Friedmann equation can be written as
$3H^{2}=-\frac{3}{4}f^{2}e^{4\pi}\dot{\pi}^{2}+\frac{3}{4}e^{2\pi}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},$ (6.10)
yielding a solution for the pressure
$p_{\pi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)\,.$ (6.11)
This shows that the universe will deviate from the quasi-Minkowskian configuration.
The background evolution of the model can be described through the Friedmann equation
$-2\dot{H}\simeq p_{\pi}=-\frac{1}{(-t)^{6}}\left(\frac{1}{2f^{2}}+\gamma\right)~{},$ (6.12)
where
$H=\frac{h_{0}}{(-t)^{5}}~{},~{}h_{0}\equiv\frac{1}{10}\left(\frac{1}{2f^{2}}+\gamma\right)~{}.$ (6.13)
Hence, the pressure and density are described by
$\rho_{\pi}=\frac{3}{4}\dot{\pi}^{4}+3H\gamma\dot{\pi}^{5}~{},~{}p_{\pi}=\frac{1}{4}\dot{\pi}^{4}-\gamma\dot{\pi}^{4}\ddot{\pi}~{},$ (6.14)
with a radiation-dominated universe solution
$a\propto t^{\frac{1}{2}}~{},~{}\dot{\pi}\propto t^{-\frac{1}{2}}\,.$ (6.15)
Hence, the universe starts from a quasi-Minkowskian state, passes through the emergent event at $t=0$, and finally gets to a standard radiation dominated universe, as in the standard cosmology.
The power spectrum for scalar and tensor perturbations of the model can be investigated by computing the second order action for linear perturbations, similarly to the standard inflationary scenario (Sec. 5) in terms of the comoving curvature perturbation $\zeta$ and then performing the field quantisation using the Mukhanov-Sasaki prescription,
[…]
where
$\frac{z_{s}^{2}}{2a^{2}}$ $=3+2\frac{\dot{\pi}^{2}(-g_{1}+3g_{2}\dot{\pi}^{2})+18H\gamma\dot{\pi}^{5}-6H^{2}}{(\gamma\dot{\pi}^{5}-2H)^{2}}\,,$ (6.17)
$\left(-\frac{z_{s}^{2}}{2a^{2}}\right)c_{s}^{2}$ $=1+h+\frac{2}{a}\frac{d}{dt}\frac{a\left(h_{X}\dot{\pi}^{2}-h-1\right)}{2H-\gamma\dot{\pi}^{5}}~{}.$ (6.18)
In the pre-emergent period $t<0$, the universe is almost static and $d\eta=dt/a\simeq dt$ and
$z_{s}^{2}=\frac{600f^{2}}{\left(1-8f^{2}\gamma\right)^{2}}(-t)^{4}~{},~{}c_{s}^{2}=\frac{4f^{2}\gamma-1}{3}\equiv c_{s0}^{2}\,,$ (6.19)
and the Mukhanov-Sasaki equation is solved in terms of Bessel functions as
$v_{k}(\eta)\propto\sqrt{k\eta}\left[c_{1}(k)J_{\frac{3}{2}}(kc_{s0}\eta)+c_{2}(k)Y_{\frac{3}{2}}(kc_{s0}\eta)\right]\,.$ (6.20)
The wavemodes of observational interests all become superhorizon in the pre-emergent period $t<0$ [118]. Hence, the dynamics of scalar perturbation have to be evaluated from past infinity to the critical time scale $\eta_{E}=t_{E}$, where the emergent solution breaks down. After $t=t_{E}$, the universe quickly turns to a radiation dominated period, and the scalar perturbation freezes.
The scalar power spectrum can be evaluated using Eq. (5.22). By using vacuum initial conditions it can be easily shown that it is scale invariant as required.
The tensor power spectrum can be computed analogously and the tensor spectral index is $n_{t}=\frac{d\ln P_{t}}{d\ln k}=2$, consistent with previous studies [125, 127].
7 Conclusions
In this review we have presented the DHOST theories – the most general scalar-tensor theories involving one scalar field, and propagating one scalar degree of freedom and two tensor ones, avoiding the generation of ghosts. These theories have been developed from the middle of the 2010s, after the observation that they do not need to involve only second-order Euler-Lagrange equations in order to satisfy this constraint, but that this can also happen when the functional coefficients in the Lagrangian satisfy certain degeneracy conditions. They include previously studied Horndeski and beyond-Horndeski theories, and while a large literature is available covering these less general models, we have not discussed their observational effects in this work, focusing only on the DHOST literature.
The research on these theories has developed into two directions in recent years. Firstly, they have been used to study the late-universe and to parameterise dark energy, incorporating previously studied models, exhibiting compatibility with the current cosmological observations and to model astrophysical objects, such as Newtonian and neutron stars as well as black holes. Secondly, they have been applied to the early universe, in the context of its early-time accelerated expansion.
This review covers both these aspects. We discuss the construction of the DHOST theories as generalisations of Horndeski and beyond-Horndeski theories, their classification, observational implications in the late universe, the use of these theories as models of dark energy and cosmological constraints, as well as inflationary model building.
Acknowledgments
This work was supported by a United Kingdom Research and Innovation (UKRI) Future Leaders Fellowship [Grant No. MR/V021974/2].